### v1

@@ III-B Peakwise Uncertainty (PU)
While MUMI is theoretically motivated, the MC-dropout requires $J$ realizations of the dropout layers to obtain $\{\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\}_{j=1}^{J}$ which is not feasible for underwater low-power applications. Here, we propose a simpler method to assess uncertainty in the classification approach tailored to the localization problem. This method counts the number of significant peaks in the output PMF, and asserts that the model is certain about the corresponding result if there is only one significant peak. Following this approach, we define a significant peak as a peak that is either the only peak or is larger than the next-largest peak by a factor $Q$. The hyperparameter $Q$ should be selected to draw a trade-off between the desired size of $\mathcal{S}$ and the minimum confidence desired in pseudo labels selected. Thus
$PU(\mathbf{x})=\begin{cases}0\quad\text{if }\hat{\mathbf{y}}\text{ has only 1 significant peak}\\ 1\quad\text{if }\hat{\mathbf{y}}\text{ has more than 1 significant peak}\end{cases}$ (10)
where $\hat{\mathbf{y}}$ is the classification network output probability vector. Throughout the paper, we use $Q=10$. The average PU indicates the percentage of samples about which the model is uncertain. Figure 4 shows that the average PU increases as the amount of mismatch increases.
@@
Figure 4: Variation of the peakwise uncertainty with the amount of environmental mismatch in (a) depth and (b) SSP.
As illustrated in the example in Fig. 5, the largest peak in the output of the pre-trained model is not necessarily aligned with the true label. The JSEA algorithm introduced in the next section, determines which peak to select.
Figure 5: An example of a predicted label in a test environment with a different SSP than the training environment. The quantization bins are $100$ m wide. Observe that the pre-trained model prediction (the label with maximum probability) is far from the ground-truth. However, there is another peak in the output closer to the ground-truth.
@@ IV-A Environmental Adaptation Using SHOT
When presented with data from an environment that is markedly different from the one on which it was trained, the network makes uncertain predictions, i.e., it is likely to be more confused between different classes, which leads to inaccurate predictions. During inference, the absence of test labels makes fully-supervised adaptation impossible. Nevertheless, encouraging the model to more confidently preserve its predictions on samples that match to the training data, helps it overcome slight mismatches. To this end, we select a self-supervising subset $\mathcal{S}$ of the samples and use the pre-trained network’s output on these samples as their pseudo-labels, for which we use a loss term $\mathsf{CE}(\mathbf{y}_{i}^{\mathsf{pseudo}},\hat{\mathbf{y}}_{i})$. Here, $\mathbf{y}_{i}^{\mathsf{pseudo}}$ is the softened label based on the output of the pre-trained classifier. However, this loss and approach can create a degenerate situation where the network only predicts classes that exist in $\mathcal{S}$ [11]. Adding a regularizing loss term to maximize the entropy of the outputs average $\bar{\mathbf{y}}=\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}\hat{\mathbf{y}}_{i}$ can prevent this situation.
At the test time, we freeze the pre-trained classifier as depicted in Fig. 1, which predicts the label $\hat{\mathbf{y}}_{i}=[\hat{y}_{i1},\hat{y}_{i2},...,\hat{y}_{iM}]$ by applying the Softmax() operation to $\boldsymbol{\omega}_{k}^{\mathsf{T}}\boldsymbol{\phi}_{i}$, i.e.,
@@
and $\boldsymbol{\omega}_{k}$ is the classification weight vector corresponding to the $k$-th class. Note that if the feature vector $\boldsymbol{\phi}_{i}$ is highly aligned with $\boldsymbol{\omega}_{k_{0}}$ for only one $k_{0}$, then the network predicts a unimodal soft label $\hat{\mathbf{y}}_{i}$, as opposed to situations where the network generates a multimodal output and is not certain about the predicted label. Inspired by this observation, we construct the self-supervising subset $\mathcal{S}$ according to
$\mathcal{S}=\{i\;:1\leq i\leq N_{test}\,\,\text{s.t.}\,\,PU(\mathbf{x}_{i})=0\}.$ (12)
@@
where, $\beta$ is a positive hyperparameter. We then employ the Adam optimizer [40] to minimize $\mathcal{L}_{\text{SHOT}}$ with a step size of $\mu_{\text{SHOT}}$. The cross-entropy term encourages the network to remain certain about its existing confident predictions, while increasing $H\left(\bar{\mathbf{y}}\right)$ prevents it from assigning all inputs to the same output by encouraging diverse outputs [11]. Although effective at rectifying errors due to small mismatches, SHOT does not directly enhance the estimates for samples in $\mathcal{S}^{c}$ and does not focus on removing the model’s confusions. To alleviate this, we now introduce another adaptation mechanism that explicitly selects pseudo-labels for the samples in $\mathcal{S}^{c}$.
@@ IV-B Joint Source-Environment Adaptation (JSEA)
While TTA methods in the machine learning literature focus on domain adaptation, they do not consider or distinguish the source that generates data and the environment where data is generated. In UWA, however, there is a clear distinction between these two factors. To improve the generalization performance of data-driven methods in UWA signal processing, a necessary step is normalization with respect to the source, which effectively yields source-invariant features and better generalization across different sources. Accordingly, array-based localization usually involves normalization by the total received signal power at the array elements. According to Fig. 6, the array SNR at the given frequency can provide a range estimate that (i) is coarse, due to the local fluctuations in the signal power level, but (ii) is robust because the general trend is that higher SNR is associated with a closer source, regardless of the environment. To exploit the distinction between the source and the environment, we implicitly estimate the test source power from a limited set of reliable samples (samples with a unimodal output PMF). We then develop a signal strength-based approach to estimate the range of more difficult samples (samples with a multimodal output PMF). This results in combining the DL-based and signal strength-based approaches in a Bayesian manner to fully use the information in an efficient manner.
@@
As a result, if $d$ is uniformly distributed over the parameter space of interest, we have $p(d|\mathbf{C})\propto p(d|\tilde{\mathbf{C}})p(d|\psi)$. The received signal power $\psi$ depends on the source power that can be estimated from $\mathcal{S}$ in (12). Note that both $p(d|\tilde{\mathbf{C}})$ and $p(d|\psi)$ can be used for localization: $p(d|\tilde{\mathbf{C}})$ provides a finer estimate based on the normalized SCM (and typically more sensitive to mismatches), while $p(d|\psi)$ provides a coarser estimate based on the received signal power that is more robust to environmental mismatches. Since both estimates use different and independent pieces of information, by exploiting both terms we can significantly improve the performance of localization algorithms in the presence of mismatches.
For each sample in $\mathcal{D}_{test}$, the pre-trained classification model generates $\hat{\mathbf{y}}_{i}=\hat{p}(d|\tilde{\mathbf{C}})$. After partitioning the samples into $\mathcal{S}$ and $\mathcal{S}^{c}$ based on their peakwise uncertainty, the final range $\hat{d}_{i}$ estimate for $\mathbf{x}_{i}\in\mathcal{S}$ is the distance corresponding to the only peak of $\hat{\mathbf{y}}_{i}$. For $\mathbf{x}_{i}\in\mathcal{S}^{c}$, let $\mathcal{P}(\hat{\mathbf{y}}_{i})$ be the set of distances corresponding to all significant peak locations of $\hat{\mathbf{y}}_{i}$. The goal is to determine a member of $\mathcal{P}(\hat{\mathbf{y}}_{i})$ as the pseudo-label for $\mathbf{x}_{i}$ by exploiting
@@
By defining the pseudo-labels for $\mathbf{x}_{i}$ as $\mathbf{y}_{i}^{\mathsf{pseudo}}\triangleq\eta(\hat{d}_{i})$, we have achieved the range estimates for all given test samples. If the model is to be used for another source in the same environment, one should also adapt the feature extractor by minimizing the JSEA loss defined by $\mathcal{L}_{\text{JSEA}}=\sum_{i=1}^{N_{test}}\mathsf{CE}(\mathbf{y}_{i}^{\mathsf{pseudo}},\hat{\mathbf{y}}_{i})$.
V Empirical Evaluation
@@ V-A Synthetic data
[…]
@@
$\text{PCL}(\zeta)$ $=\frac{100}{N_{test}}\sum_{i=1}^{N_{test}}\mathbf{1}_{|d_{i}-\hat{d}_{i}|\leq\zeta d_{i}},$
MAE $=\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}|d_{i}-\hat{d}_{i}|,$
@@
  • M-MFP (mismatched MFP): The MFP approach tuned to the training environment. This method, in contrast with data-driven methods, has access to the training data.
  • CNN-c: A CNN-based model for classification.
@@
Figure 8 shows the superiority of the JSEA-c and CNN-c methods over M-MFP when the training and testing environments have a depth mismatch of $\Delta D=4$ m (i.e, the ocean depth in the test environment is $220.5$ m), and there is a mismatch in the noise statistics of the test data compared to the training data as well. The proposed JSEA-c outperforms other methods including SHOT, however, its improvement highly depends on the success of the pre-trained model in generating enough certain labels. Although, JSEA-r can successfully improve the regression results as seen in Fig. 8, we will not include the regression models in the following experiments due to their inferior performance. In this scenario, the M-MFP shows inferior performance with respect to classification methods, in the presence of depth mismatch.
Figure 9 shows the performance comparison of the methods with variation in the degree of SSP mismatch. In the current scenario, mismatch in SSP gradient is assessed, by varying the gradient difference $\Delta c$ between training and test data. As seen in Fig. 9, the M-MFP outperforms the DL models; however, it benefits from using training data during inference, which is not available to the other methods. While SHOT can improve the CNN-c results, JSEA-c has the best performance among the DL methods.
@@
Next we consider the depth mismatch and evaluate the performance of the algorithms for $\Delta D\in\{0,2,4,6\}$ m/s. As seen in Fig. 10, when the test and training environments match (i.e., $\Delta D=0$), the MFP (in this case, the M-MFP is not mismatched, hence it is the same as O-MFP) performs the best. However, depth mismatches lead to the M-MFP’s performance deterioration, whereas the performance of the CNN-based methods is more robust to this mismatch. Thus, using CNN methods can impart some robustness to localization performance under depth mismatch.
(a) PCL(10%) under mismatch in SSP.
(b) MAE under mismatch in SSP.
@@ V-B SWellEx-96 Data
To obtain the Fourier transform and the corresponding SCM $\mathbf{C}$ used for localization, we use a $3$-second-long signal divided into $P=5$ segments, each of length $1$ second and a $50\%$ overlap, tapered by a Kaiser window with $\beta=9.24$. We obtain the true source range by linear interpolation between the GPS recordings provided.
Figure 11: The ship track for the SWellEx-96 S5 event [41]. We aim to estimate the distance of the ship to the VLA receiver.
Figure 12 shows the ranging results of M-MFP, CNN-c, and JSEA-c on the real-world narrow-band data at frequency $f=109$ Hz. The results in Fig. 12(a) suggest that M-MFP performs poorly in this scenario, as it leans towards overestimating the ranges for the samples around $t=60$ min which are closest to the VLA and have a relatively higher SNR. This problem is absent in the CNN-c approach as observed in Fig. 12(b). Nevertheless, both M-MFP and CNN-c tend to deviate significantly from the true ranges in the low SNR regions, i.e., larger source ranges ($t\leq 40$ min). Whereas this problem is still present in JSEA-c, it is considerably less severe, yielding an overall better MAE and PCL, as shown in Table I.
@@
TABLE I: Localization performance for SWellEx-96 data.
@@
VI Conclusion

### v2

@@ III-B Peakwise Uncertainty (PU)
While MUMI is theoretically motivated, the MC-dropout requires $J$ realizations of the dropout layers to obtain $\{\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\}_{j=1}^{J}$ which is not feasible for underwater low-power applications. Here, we propose a simpler method to assess uncertainty in the classification approach tailored to the localization problem. This method counts the number of significant peaks in the output PMF, and is based on the intuition that the model is certain about the corresponding result if there is only one significant peak, whereas in cases where noise or environmental mismatch dominate, there would not be one significant peak only.Following this approach, we define a “significant peak” as a peak that is either the only one or is larger than the next-largest peak by a factor of at least $Q$. The hyperparameter $Q$ should be selected to draw a trade-off between the desired size of $\mathcal{S}$ and the minimum confidence desired in pseudo labels selected. Thus
[…]
where $\hat{\mathbf{y}}$ is the classification network output probability vector. Throughout the paper, we use $Q=10$. The average PU indicates the percentage of samples about which the model is uncertain and is defined as
$APU[\%]=100\frac{\sum_{i=1}^{N}PU(\mathbf{x}_{i})}{N}.$
Fig. 4 shows that the APU increases as the amount of mismatch increases.
@@
Figure 4: Variation of the average peakwise uncertainty with the amount of environmental mismatch in (a) depth and (b) SSP.
Peak counting is not commonly used for uncertainty quantification. Nevertheless, APU shows a similar variation as MUMI in the presence of mismatches. This motivates interpreting PU as an implied uncertainty. As illustrated in the example in Fig. 5, the largest peak in the output of the pre-trained model is not necessarily aligned with the true label. The JSEA algorithm introduced in the next section, determines which peak to select.
Figure 5: An example of a predicted label in a test environment with a different SSP than the training environment. The quantization bins are $100$ m wide. Observe that the pre-trained model prediction (the label with maximum probability) is far from the ground-truth. However, there is another peak in the output closer to the ground-truth. After JSEA-based adaptation, the output shows one clear peak at the true label, showing how JSEA corrects and adapts the network to give correct labels in the presence of environmental mismatch.
@@ IV-A Environmental Adaptation Using SHOT
When presented with data from an environment that is markedly different from the one on which it was trained, the network makes uncertain predictions, i.e., it is likely to be more confused between different classes, which leads to inaccurate predictions. During inference, the absence of test labels makes fully-supervised adaptation impossible. Nevertheless, encouraging the model to more confidently preserve its predictions on samples that match to the training data, helps it overcome slight mismatches. To this end, we select a self-supervising subset $\mathcal{S}$ of the samples and use the pre-trained network’s output on these samples as their pseudo-labels, for which we use a loss term $\mathsf{CE}(\mathbf{y}_{i}^{\mathsf{pseudo}},\hat{\mathbf{y}}_{i})$. Here, $\mathbf{y}_{i}^{\mathsf{pseudo}}$ is the softened label based on the output of the pre-trained classifier. However, this loss and approach can create a degenerate situation where the network only predicts classes that exist in $\mathcal{S}$ [17]. Adding a regularizing loss term to maximize the entropy of the outputs average $\bar{\mathbf{y}}=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\hat{\mathbf{y}}_{i}$ can prevent this situation.
At test time, we freeze the pre-trained classifier as depicted in Fig. 1, which predicts the label $\hat{\mathbf{y}}_{i}=[\hat{y}_{i1},\hat{y}_{i2},...,\hat{y}_{iM}]$ by applying the Softmax() operation to $\boldsymbol{\omega}_{k}^{\mathsf{T}}\boldsymbol{\phi}_{i}$, defined as,
@@
where $\boldsymbol{\omega}_{k}$ is the classification weight vector corresponding to the $k$-th class. Note that if the feature vector $\boldsymbol{\phi}_{i}$ is highly aligned with $\boldsymbol{\omega}_{k_{0}}$ for only one $k_{0}$, then the network predicts a unimodal soft label $\hat{\mathbf{y}}_{i}$, as opposed to situations where the network generates a multimodal output and is not certain about the predicted label. Inspired by this observation, we construct the self-supervising subset $\mathcal{S}$ according to
$\mathcal{S}=\{i\;:1\leq i\leq N_{\text{test}}\,\,\text{s.t.}\,\,PU(\mathbf{x}_{i})=0\}.$ (12)
@@
where $\beta$ is a positive hyperparameter. We then employ the Adam optimizer [45] to minimize $\mathcal{L}_{\text{SHOT}}$ with a step size of $\mu_{\text{SHOT}}$ to update the CNN feature extractor. The cross-entropy term encourages the network to remain certain about its existing confident predictions, while increasing $H\left(\bar{\mathbf{y}}\right)$ prevents it from assigning all inputs to the same output by encouraging diverse outputs [17]. Fig. 6 provides a flowchart for adapting the model using the SHOT algorithm.
Figure 6: A flowchart for the SHOT approach for adapting the localization model. There are three major steps in SHOT: First, the CNN model estimates the labels; then the set of certain samples $\mathcal{S}$ is determined. For the samples in $\mathcal{S}$, the CNN estimates are considered the final range estimates (pseudo-labels). Then, the CNN is retrained using the pseudo-labeled samples in $\mathcal{S}$. Finally, for each of the uncertain samples in $\mathcal{S}^{c}$, the final range estimate is obtained from the retrained CNN.
Although effective at rectifying errors due to small mismatches, SHOT does not directly enhance the estimates for samples in $\mathcal{S}^{c}$ and does not focus on removing the model’s confusions. To alleviate this, we now introduce another adaptation mechanism that explicitly selects pseudo-labels for the samples in $\mathcal{S}^{c}$.
@@ IV-B Joint Source-Environment Adaptation (JSEA)
While TTA methods in the machine learning literature focus on domain adaptation, they do not consider or distinguish the source that generates data and the environment where data is generated. In UWA, however, there is a clear distinction between these two factors. To improve the generalization performance of data-driven methods in UWA signal processing, a necessary step is normalization with respect to the source, which effectively yields source powerinvariant features and better generalization across different sources. Accordingly, array-based localization usually involves normalization by the total received signal power at the array elements. However, by doing this, one loses valuable information on the source power that could be leveraged to improve localization. For example, as shown in Fig. 7, the array SNR at the given frequency can provide a range estimate that (i) is coarse, due to the local fluctuations in the signal power level, but (ii) is robust because the general trend is that higher SNR is associated with a closer source, regardless of the environment. To exploit the distinction between the source and the environment, we implicitly estimate the test source power from a limited set of reliable samples (samples with a unimodal output PMF). We then develop a signal strength-based approach to estimate the range of more difficult samples (samples with a multimodal output PMF). This results in combining the DL-based and signal strength-based approaches in a Bayesian manner to fully use the information in an efficient manner.
@@
[…]
@@
By defining the pseudo-labels for $\mathbf{x}_{i}$ as $\mathbf{y}_{i}^{\mathsf{pseudo}}\triangleq\eta(\hat{d}_{i})$, we have achieved the range estimates for all given test samples. If the model is to be used for another source in the same environment, one should also adapt the feature extractor by minimizing the JSEA loss defined by $\mathcal{L}_{\text{JSEA}}=\sum_{i=1}^{N_{\text{test}}}\mathsf{CE}(\mathbf{y}_{i}^{\mathsf{pseudo}},\hat{\mathbf{y}}_{i})$. Fig. 8 provides a flowchart for the JSEA algorithm.
While SHOT defines the pseudo-labels only for the samples in $\mathcal{S}$, JSEA defines pseudo-labels for all test samples. In both SHOT and JSEA, pseudo-labels for certain samples are the final range estimates. However, to obtain the final range estimates for uncertain samples, SHOT relies on re-training the network for a few iterations using the pseudo-labeled certain samples, while JSEA relies on the peak selection process based on the received signal power.
Figure 8: A flowchart for JSEA. There are two major steps in JSEA: First, the CNN model estimates the labels; then the set of certain samples $\mathcal{S}$ is determined. For the samples in $\mathcal{S}$, the CNN estimates are considered the final range estimates. Finally, for each of the uncertain samples in $\mathcal{S}^{c}$, the final range estimate is selected from the set of peak locations, according to (17).
V Evaluation
@@ V-A Synthetic data
$\text{SNR [dB]}=10\log_{10}\frac{\sum_{i=1}^{N_{\text{test}}}\sum_{l=1}^{L}|r_{i,l}|^{2}}{N_{\text{test}}L\psi_{w}},$
@@
$\text{PCL}(\zeta)$ $=\frac{100}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\mathbf{1}_{|d_{i}-\hat{d}_{i}|\leq\zeta d_{i}},$
MAE $=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}|d_{i}-\hat{d}_{i}|,$
@@
  • M-MFP (mismatched MFP): The MFP approach tuned to the training environment. This method is assumed to have access to the training data.
  • CNN-c: A CNN-based model for classification, trained once on the training set, and not adapted to account for mismatch.
@@
Fig. 10 shows the superiority of the JSEA-c and CNN-based classification methods over M-MFP when the training and testing environments have a depth mismatch of $\Delta D=4$ m (i.e, the ocean depth in the test environment is $220.5$ m), and there is a mismatch in the noise statistics of the test data compared to the training data as well. The proposed JSEA-c outperforms other methods including SHOT, however, its improvement highly depends on the success of the pre-trained model in generating enough certain labels. Although, JSEA-r can successfully improve the regression results as seen in Fig. 10, we will not include the regression models in the following experiments due to their inferior performance compared to the classifier-based approaches. In this scenario, the M-MFP shows inferior performance with respect to classification methods, in the presence of depth mismatch.
@@
Fig. 11 shows the performance comparison of the methods with variation in the degree of SSP mismatch. In the current scenario, mismatch in SSP gradient is assessed, by varying the gradient difference $\Delta c$ between training and test data. As seen in Fig. 11, the M-MFP outperforms the DL models; however, it benefits from using training data during inference, which is not available to the other methods. While SHOT can improve the CNN-c results, JSEA-c has the best performance among the DL methods.
@@
Next we consider the depth mismatch and evaluate the performance of the algorithms for $\Delta D\in\{0,2,4,6\}$ m/s. As seen in Fig. 12, when the test and training environments match (i.e., $\Delta D=0$), the MFP (in this case, the M-MFP is not mismatched, hence it is the same as O-MFP) performs the best. However, depth mismatches lead to the M-MFP’s performance deteriorating, whereas the performance of the CNN-based methods is more robust to this mismatch. Thus, using CNN methods can impart some robustness to localization performance under depth mismatch.
(a) PCL(10%) under mismatch in depth.
(b) MAE under mismatch in depth.
@@
To further investigate the effect of environmental mismatches, we evaluate the performance against mismatches in bathymetric shape/fluctuations, as shown in Fig. 13. While the training environment includes a flat bathymetry, the test environment contains a seamount. Fig. 14 shows the MAE and PCL results versus the seamount height. The results in Fig. 14 verify the superior performance of the JSEA-c in terms of both MAE and PCL for this scenario.
(a) Bathymetry of the training environment, inspired by the SWellEx-96 environment.
(b) Mismatched bathymetry used for testing. There is a seamount structure at 1 km distance from the source, with a height of $h$ meters.
Figure 13: Bathymetry mismatch used to evaluate the performance of the JSEA method.
(a) PCL(10%) under mismatch in bathymetry.
(b) MAE under mismatch in bathymetry.
Figure 14: Performance against variation in bathymetry at an SNR of $15$ dB.
Mismatch in bottom sediment type is another type of environmental mismatch that commonly arises in localization problems. Each sediment type varies in density, sound speed, and attenuation coefficient, which affect the acoustic wave propagation through bottom interactions. For the sake of demonstrating robustness to mismatch, we consider only fluid sediment types here, and do not consider shear speed or attenuation. To evaluate the performance against mismatches in sediment type, we generate training data according to the sediment type provided in Fig.13(a) and evaluate the performance on test data generated in environments with sediment types defined according to Table I [48]. Fig. 15 shows the MAE and PCL results versus the sediment type. These results verify the superior performance of the JSEA-c in terms of both MAE and PCL for this scenario.
TABLE I: Different sediment types [48]
training sediment clay silt sand gravel moraine
density $\rho\;\;[g/cm^{3}]$ 1.76 1.5 1.7 1.9 2 2.1
minimum sound speed $c_{\text{min}}[m/s]$ 1572.37 1500 1575 1650 1800 1950
maximum sound speed $c_{\text{max}}[m/s]$ 1593.02 1520 1595 1670 1820 1970
attenuation $\alpha$ [dB/ (km Hz)] 0.2 0.2 1 0.8 0.6 0.4
(a) PCL(10%) under mismatch in sediment type.
(b) MAE under mismatch in sediment type.
Figure 15: Performance against mismatch with various sediment types at an SNR of $15$ dB.
@@ V-B SWellEx-96 Data
To obtain the Fourier transform and the corresponding SCM $\mathbf{C}$ used for localization, we use a $3$-second-long signal divided into $P=5$ segments, each of length $1$ second and a $50\%$ overlap, tapered by a Kaiser window with $\beta=9.24$. We obtain the true source range by linear interpolation between the GPS recordings provided assuming the ship followed a straight-line path between GPS updates.
Figure 16: The ship track for the SWellEx-96 S5 event [46]. The localization problem considered involves estimating the distance of the ship to the VLA receiver.
Fig. 17 shows the ranging results of M-MFP, CNN-c, and JSEA-c on the real-world narrowband data at frequency $f=109$ Hz. The results in Fig. 17(a) suggest that M-MFP performs poorly in this scenario, as it leans towards overestimating the ranges for the samples around $t=60$ min which are closest to the VLA and have a relatively higher SNR. This problem is absent in the CNN-c approach as observed in Fig. 17(b). Nevertheless, both M-MFP and CNN-c tend to deviate significantly from the true ranges in the low SNR regions, i.e., larger source ranges ($t\leq 40$ min). While this problem is still present in JSEA-c, it is considerably less severe, yielding an overall better MAE and PCL, as shown in Table II.
@@
TABLE II: Localization performance for SWellEx-96 data.
@@
VI Discussion, Extensions, And Limitations of JSEA
Whereas the empirical evaluations in the previous section verify the efficacy of JSEA in several different scenarios, it is a sophisticated method with some inherent limitations. This section addresses hyperparameter selection, JSEA limitations and possible algorithmic extensions, and complexity comparison with other methods.
VI-A Hyperparameter Selection
Successful operation of JSEA relies on reasonable selection of the hyper-parameters. The hyper-parameters $\sigma$ and $B$ are related to converting the regression problem to a classification one, and can be selected by evaluating the performance on a validation set, and adjusting their values according to the following guidelines. Moreover, the adaptation hyper-parameters $\delta$ and $Q$ can be determined based on the desired performance and any prior knowledge on the degree of the mismatch between training and test environments, according to the following discussion. In the following, we explain the effect of each of these four hyper-parameters.
  • $\sigma$: A large $\sigma$ leads to ground-truth labels that are heavy-tailed and closer to a uniform distribution and cannot indicate the distinction between different classes clearly. This in turn slows the training due to weak supervisory signal and can lead to an inferior accuracy. On the other hand, a small $\sigma$ leads to ground-truth labels that are light-tailed and closer to a point mass (spike), rendering labels one-hot-coded, which can exacerbate the overfitting of the model by removing the distance information between different classes. Hence the $\sigma$ value is chosen to balance the tradeoff between these two factors.
  • $B$: A larger $B$ leads to lower precision (a coarser quantization) in localization which bounds the best localization performance possible, but also yields a less complex model with fewer output classes which makes classification easier and enhances performance. So in practice, $B$ should be determined according to this tradeoff between the precision required and the desired model complexity. Based on this tradeoff, and also following Chen and Schmidt [1], who used a similar CNN-based architecture for source ranging using the SWellEx-96 dataset, we select $B=100$ m.
  • $Q$: The hyperparameter $Q$ adjusts the sensitivity of the algorithm to mismatches; the larger $Q$ is, the more sensitive JSEA becomes to mismatches. As such, $Q$ should be selected to draw a trade-off between the desired size of $\mathcal{S}$ and the minimum confidence desired in pseudo-labels selected. Determining this trade-off is naturally tied to $\sigma$ and $\delta$. We empirically observed that $Q=10$ leads to a reasonable trade-off.
  • $\delta$: This hyper-parameter should be small enough to be able to distinguish among the peaks to rectify the labels of uncertain samples, i.e.,
    $\forall j\in\mathcal{S}^{c}:\;\;\forall d_{1},d_{2}\in\mathcal{P}(\hat{\mathbf{y}}_{j}),\;\;\text{if}\;\;d_{1}\neq d_{2},\;\;\mathcal{S}_{\delta}(d_{1})\neq\mathcal{S}_{\delta}(d_{2}),$
    and it should be large enough that
    $\forall j\in\mathcal{S}^{c}:\;\;\forall d\in\mathcal{P}(\hat{\mathbf{y}}_{j}),\;\;\mathcal{S}_{\delta}(d)\neq\phi.$
    If such a $\delta$ cannot be found for a sample $j$, one can either accept the largest output peak as the final estimate, or declare the sample as an uncertain one that cannot be rectified using JSEA. We have used $\delta=500$ m for all test samples as this value satisfied these conditions.
VI-B Complexity Comparison
Underwater applications typically require power- and electronic storage-efficient algorithmic designs. Here, we provide a complexity comparison between different methods that were discussed in the paper. Although the calculations are not based on hardware-optimized algorithms, they provide a reasonable baseline for comparison. The complexity analysis is provided in Appendix A. Table III summarizes the complexity and run-time comparison between different methods, when implemented on an NVIDIA RTX 6000 Ada Generation GPU device using CUDA 12.2.
TABLE III: Memory and Computational Complexity Comparison
Model Computational Complexity Memory Complexity Run Time $[ms]$
MFP $O(N_{\text{tr}}L^{2})$ $O(N_{\text{tr}}L)$ $0.4666$
CNN-c $O(N_{\phi}L^{2}+MN_{\phi})$ $O(N_{\phi}L^{2}+MN_{\phi})$ $0.0014$
SHOT $O(N_{\phi}L^{2}+MN_{\phi}+N_{\text{itr}}N_{\phi}L^{2})$ $O(N_{\text{test}}L+N_{\phi}L^{2}+MN_{\phi})$ $1.5225$
JSEA-c $O(N_{\phi}L^{2}+MN_{\phi}+N_{\text{test}})$ $O(N_{\text{test}}L+N_{\phi}L^{2}+MN_{\phi})$ $0.2353$
VI-C Extensions And Limitations
The proposed JSEA mechanism relies on access to a sufficient number of test samples as well as the ability of the CNN method to classify a sufficient number of them with certainty. Determining the minimum number of certain samples required for successful operation of JSEA is left for future research. Furthermore, variations of JSEA can be developed that use other assumptions for $p(d|\psi)$ in (15) and correspondingly may lead to more efficient methods for source power estimation from the samples in $\mathcal{S}$.
In the presence of more than one acoustic source at the intended frequency, MFP and CNN-based methods may fail in localization. In such scenarios, the JSEA method might show spurious peaks even in an environment that closely matches the training environment, i.e., JSEA reveals a mismatch in the situation (multi-source situation during inference versus the single-source situation during the training). Nevertheless, JSEA is not able to rectify such errors, as it requires samples from the same source at different ranges to implicitly estimate the source power.
The simulations in this paper verify the efficacy of JSEA for narrowband sources. For broadband source localization, there are several approaches to exploit the information contained in different frequency bands. One approach is using the SCMs obtained for multiple frequencies as inputs of size $2N_{F}\times L\times L$ to the CNN model, where $N_{F}$ is the number of frequencies. In this scenario, JSEA can be used in the same manner it is used the narrowband case. However, another approach is to train a different sub-model for each frequency and then aggregate their estimates to obtain the source range. In this scenario, we can either apply JSEA to each sub-model first and then aggregate the results, or first aggregate the results (e.g., by averaging the output PMFs from all sub-models) and then use JSEA as in the narrowband scenario. The efficacy of this approach remains to be explored in future work.
VII Conclusion